Project

General

Profile

NDOS CommissionMtng 20121112
Attending: Jon, Jan, Susan, Nick, Kurt, Junting, Venktesh, Peter, Jaraslov, Xuebing, Leon, Denis

DataCheck - Nearline web interface offline. Jarek disable to avoid flooding of emails during DDF work.
- This weekend empty events seens in 14301-03 (subrun 20, 1/3)
- DCSReports: has time stamp has changed: we're not aware of any changes to the table
- Run 14300: Two shut off on 2-2 40, 60, 0 tecc.rval (What does zero mean here?)

DDF - Denis was working on DDF
- DSO worked fine
- DCS app were crashing with exceptions : Denis figured out most of the issues
- Rolled back
- Ron and Denis maybe have identified the problem.
- Continue Today.
- Lots of rebooting of DCM. Detector came back noisy. Force reprogramming needed.

Steven - Foulkes IOC sync. Has fix installing today

Junting - Showed more on multiplexing issue. Still inconclusive as to if multiple pixels are being sampled where one is expected. (NOVA-doc-8197-v1)

Venktesh - Pulser test last week. Spacer frame could see the pulse. Greg working on permenant mount.
- Greg will ready this week with LED mount and wants to test fit
- Multichannel ready at the end of the week

APD Teststand - Joao teststand DAQ (Adrew is gone this week)
- 10 FEB assembled
- 20 FEB later today
- Leons blood pressure is up
- APDs Ship out tomorrow so we'll see them on Wednesday with Leon.

Jon - Big changes to the software with DAQChannelMaps (Roll out to test stand soon)